# A stochastic Gordon-Loeb model for optimal cybersecurity investment under clustered attacks

### v1

@@ 6 Conclusions
In this work, we introduce a dynamic and stochastic extension of the Gordon–Loeb model [3] for optimal cybersecurity investment, incorporating temporally clustered cyberattacks via a Hawkes process. Our modeling framework captures the empirically observed phenomenon of attack bursts, thus offering a more realistic representation of the current cyber-risk environment. We formulate the cybersecurity investment decision problem as a two-dimensional stochastic optimal control problem, maximizing the expected net benefit of cybersecurity investments. We allow for adaptive investment policies that respond in real-time to the arrival of cyberattacks.
Our numerical results demonstrate that the optimal cybersecurity investment policy consistently outperforms both static benchmarks and Poisson-based models that ignore clustering. In particular, even when Poisson models are calibrated to match the expected attack frequency, they fail to capture the implications of attack clustering on investment timing and magnitude, thus leading to suboptimal investment decisions. Our findings indicate that the optimal dynamic strategy is able to react promptly to attack clusters, offering substantial improvements in expected net benefit in high-risk scenarios. Overall, our results underscore the importance of accounting for dynamic and stochastic threat patterns in cybersecurity planning. The proposed framework supports risk managers and policymakers in designing responsive cybersecurity investment strategies tailored to the evolving cyber-risk landscape.
Future research directions include the empirical calibration to sector-specific cyber incident data, the consideration of risk-aversion with respect to losses resulting from cyberattacks, and the integration of cyber-insurance as a complementary tool for risk mitigation (see [34, 35, 28, 7, 36, 27] for some recent studies in this direction). Our framework can also be applied from the viewpoint of an insurance firm which provides insurance against losses due to cyberattacks, thus laying the foundations for the development of Cramér-Lundberg-type models (see, e.g., [37]) for cyber-insurance. Finally, our modeling setup can also be extended to multivariate Hawkes processes (as considered in [38], or in the more general versions of [39, 40]) to differentiate among multiple types of cyberattacks (see [41] for a multivariate generalization of the static Gordon-Loeb model).
@@ Acknowledgements
This work benefited from financial support from the Europlace Institute of Finance, Fondazione Gini, Unione Matematica Italiana and from the European Union-Next Generation EU - PRIN research project n. 2022BE-MMLZ. The authors thank Alessandro Calvia, Salvatore Federico, Luca Grosset, Eleonora Losiouk, Daniele Marazzina, Andrea Pallavicini, Andrea Perchiazzo, Mario Putti, Edit Rroji for valuable discussions on the topic of this work.

### v2

@@ 7 Conclusions
In this work, we introduce a dynamic and stochastic extension of the Gordon-Loeb model [24] for optimal cybersecurity investment, incorporating temporally clustered cyberattacks via a Hawkes process. Our modeling framework captures the empirically observed phenomenon of attack bursts, thus offering a more realistic representation of the current cyber-risk environment. We formulate the cybersecurity investment decision problem as a two-dimensional stochastic optimal control problem, maximizing the expected net benefit of cybersecurity investments. We allow for adaptive investment policies that respond in real-time to the arrival of cyberattacks.
Our numerical results demonstrate that the optimal cybersecurity investment policy consistently outperforms both static benchmarks and Poisson-based models that ignore clustering. In particular, even when Poisson models are calibrated to match the expected attack frequency, they fail to capture the implications of attack clustering on investment timing and magnitude, thus leading to suboptimal investment decisions. Our findings indicate that the optimal dynamic strategy is able to react promptly to attack clusters, offering substantial improvements in expected net benefit in high-risk scenarios. From an actuarial perspective, we interpret cybersecurity investment as a form of self-insurance, showing that improved prevention lowers both expected losses and loss variability, which in turn reduces cyber-insurance premia under standard premium principles. Overall, our results underscore the importance of accounting for dynamic and stochastic threat patterns in cybersecurity planning. The proposed framework supports risk managers, insurers and policymakers in designing responsive cybersecurity investment strategies and cyber-insurance pricing schemes that are consistent with the evolving cyber-risk landscape.
Future research directions include the empirical calibration to sector-specific cyber incident data, the consideration of risk-aversion with respect to losses resulting from cyberattacks, and the integration of cyber-insurance as a complementary tool for risk mitigation (see [1, 18, 43, 45, 49, 55, 54] for some recent studies in this direction). Our framework can also be applied from the viewpoint of an insurance firm which provides insurance against losses due to cyberattacks, thus laying the foundations for the development of Cramér-Lundberg-type models (see, e.g., [46]) for cyber-insurance. Finally, our modeling setup can also be extended to multivariate Hawkes processes (as considered in [19], or in the more general versions of [7, 8]) to differentiate among multiple types of cyberattacks (see [5] for a multivariate generalization of the static Gordon-Loeb model).
@@ Acknowledgements
The authors are thankful to Alessandro Calvia, Salvatore Federico, Luca Grosset, Eleonora Losiouk, Daniele Marazzina, Andrea Pallavicini, Andrea Perchiazzo, Mario Putti, Edit Rroji for valuable discussions on the topic of this work. Constructive comments by three anonymous reviewers and an associate editor are gratefully acknowledged.
Competing Interests
All authors declare they do not have competing interests.
Data Availability Statement
Data availability is not applicable to this article as no new data were created or analysed in this study. If needed, the authors agree to make available the code used for the numerical results illustrated in Section 5.
Funding
This work was supported by the Europlace Institute of Finance; Fondazione Gini; Unione Matematica Italiana and European Union-Next Generation EU [PRIN-2022BE-MMLZ]; and the Chair Stress Test, RISK Management and Financial Steering, led by the French Ecole Polytechnique and its Foundation and sponsored by BNP Paribas. The funders had no role in study design, decision to publish, or preparation of the manuscript.